# Classification of fractional, singular Yamabe metrics on a twice-punctured sphere I

### v1

@@ 1. Introduction
where $\in(0,\frac{n}{2}), s\mathbb{P}_{s}$ is the family of fractional Paneitz operators constructed by Graham and Zworski in [graham-zworski] (see also [chang-gonzalez]) and $c_{n,s}>0$ is the normalization constant defined as
@@
We are mostly interested in this equation on subdomains of the sphere, i.e. $\Omega:=\mathbf{S}^{n}\setminus\Lambda$ with $\Lambda\subset\mathbf{S}^{n}$ a closed set. We denote the usual round metric as $g_{\circ}$. In this locally conformally flat setting, $\mathbb{P}_{s}$ has the explicit form (originally given by Branson in [Branson_funct_det]) given by
@@
Later, we will see a similar expansion for $\mathbb{P}_{s}$ in the cylindrical setting, which was first derived in [Azahara2018].
@@
$\hat{c}_{n,s}=2^{2s-n}{2}\left(\frac{\Gamma\left(\frac{1}{2}\left(\frac{n}{2}+s\right)\right)}{\Gamma\left(\frac{1}{2}\left(\frac{n}{2}-s\right)\right)}\right)^{\frac{2s-n}{2s}}\left(\frac{\Gamma\left(\frac{n}{2}-s\right)}{\Gamma\left(\frac{n}{2}+s\right)}\right)^{\frac{2s-n}{4s}}.$
Here we remark that $\hat{c}_{n,s}>1$ (cf. [Azahara2018, Proposition 6.1]).
@@
In previous work, DelaTorre, del Pino, González and Wei [MR3694655] constructed a family of solutions on the twice-punctured sphere. These solutions are, after an appropriate conformal dilation, all rotationally symmetric. They generalize the well-known Delaunay metrics in the scalar curvature setting, which were found by Schoen [MR994021], and we call them Delaunay metrics as well. A Delaunay metric is uniquely determined by its fundamental period, or equivalently by its necksize. More precisely, the fractional Delaunay family can be equivalently characterized as the set of even, positive, periodic solutions $v_{\varepsilon}:\mathbf{R}\to(0,\infty)$ of a nonlocal initial value problem (see ($\widehat{\mathcal{Q}}_{n,s}$)) that fixes the necksize parameter $\varepsilon=\min_{t\in\mathbf{R}}v_{\varepsilon}(t).$ For each necksize $\varepsilon\in(0,1]$, the Euclidean conformal factor
@@
$g_{\varepsilon}=U_{\varepsilon}^{\frac{4}{n-2s}}g_{\circ},\qquad u_{\varepsilon}=U_{\varepsilon}\cdot u_{\circ}.$
@@
Let $n\in\mathbf{N}$ with $n\geq 3$. There exists $\delta>0$ such that if $s\in(1-\delta,1)$ and $u\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$ is a solution to ($\overline{\mathcal{Q}}_{n,s}$), then there exists $\varepsilon\in(0,1]$ and $R>0$ such that
$u=u_{\varepsilon,R},\qquad u_{\varepsilon,R}(x)=R^{\frac{2s-n}{2}}u_{\varepsilon}\left(\frac{x}{R}\right),$
where $u_{\varepsilon}$ is the Euclidean Delaunay solution of necksize $\varepsilon$.
@@
Let $n\in\mathbf{N}$ with $n\geq 3$. There exists $\delta>0$ such that if $s\in(1-\delta,1)$ then $\mathcal{M}_{\{p,q\}}=(0,1]$ for each pair of points with $p\neq q$.
We use the same technique as in the proof of Theorem 1 to prove that all Delaunay solutions are nondegenerate. This condition is easiest to define in the cylindrical gauge.
Definition 3.
Let $v:\mathbf{R}\times\mathbf{S}^{n-1}\rightarrow(0,\infty)$ solve ($\widehat{\mathcal{Q}}_{n,s}$) and let
@@
be the associated linearized operator. We say that $v$ is nondegenerate if $w\in L^{2}(\mathbf{R}\times\mathbf{S}^{n-1})$ and $\mathbb{L}_{v}(w)=0$ implies $w\equiv 0$.
@@
Let $n\in\mathbf{N}$ with $n\geq 3$. There exists $\delta>0$ such that if $s\in(1-\delta,1)$ and $v:\mathbf{R}\times\mathbf{S}^{n-1}\rightarrow(0,\infty)$ satisfies ($\widehat{\mathcal{Q}}_{n,s}$) then $v$ is nondegenerate.
In the scalar curvature setting, the conformal factor satisfies a second order ODE, and so the uniqueness follows very quickly from phase-plane analysis. However, as demonstrated in [Azahara2018], one cannot apply these techniques directly in the nonlocal setting. Nonetheless, we can prove uniqueness if $s$ is close to, but less than one.
Our proof of uniqueness relies on two new ideas. In § 3.2, we use the fact that the spherical solution has Morse index one to derive a sharp a priori upper bound for a solution to the $Q$-curvature equation of order $s$ on a twice-punctured sphere, with the spherical solution giving us a universal upper bound. This allows us to associate a Delaunay solution with the same $L^{\infty}$-norm to any solution on the twice-punctured sphere. We prove Theorem 1 by contradiction, assuming there exists a sequence $s_{k}\nearrow 1$ and solutions $v_{k}$ that are not Delaunay solutions. Letting $\overline{v}_{k}$ be the matched Delaunay solution, in § 4.1 we prove that a rescaling of the difference $v_{k}-\overline{v}_{k}$ converges locally to a solution of a homogeneous second order ODE. In § 4.2 we extend this local convergence to global convergence. This limit must be zero by the conditions of the ODE, which contradicts the normalization of our rescaling. The passage from local convergence to global convergence relies on a sharp estimate for the integral kernel and the solution operator of the nonlocal equation, derived in [MR3694655] and described below in § 2.2. We expect this strategy to be robust and to have applications in other fractional geometric problems, and indeed with adapt our technique to prove Theorem 4 in § 5. We conclude this paper with a brief outlook in § 6. First, in § 6.1, we discuss work in progress [our_followup] in which we extend the classification to $s\in(1,1+\delta)$, where once again $\delta>0$ is sufficiently small. Next, in § 6.2, we describe some possible applications of our classification result, including refined asymptotics of solutions with isolated singularities and a general compactness result.
@@ 2. Qualitative properties of Delaunay solutions
In this section, we first describe the reduction to a one-dimensional problem. Next, we describe the Delaunay solutions. Finally, we recall that Dávila, Del Pino and Sire [DDS] (see also [Andrade_Wei_Ye]) proved that the standard bubble has Morse index one.
@@ 2.1. Rotational symmetry
One can now apply [MR3366748] to see that $u$ is rotationally invariant. Alternatively, one can argue as follows. Ngô and Ye [MR4438901] prove that, because the singular set $\{0\}$ has capacity 0, the function $u$ is $s^{\prime}$-superharmonic for each $0<s^{\prime}<s$. One can now use the method of moving planes to show that $u$ is radial. One can also use the method of moving planes in integral form (see [MR3626036] and [MR2200258]) to show the radial symmetry directly.
@@ 2.2. Existence of Delaunay solutions
The construction in [MR3694655] starts with the Ansatz that the solution $u$ has the form
@@
where $\widetilde{v}:(0,\infty)\rightarrow(0,\infty)$ is a positive smooth function. In fact, one can use the a priori bounds of [MR3366748] to show that $\widetilde{v}$ is bounded, and so the integral operators written below are convergent. This change of variables transforms ($\overline{\mathcal{Q}}_{n,s}$) into
@@
is a positive constant; in fact, our normalizations imply $A=c_{n,s}$ (see [Azahara2018, Proposition 2.7]). Finally, we complete the Emden-Fowler change of variables by letting $t=-\ln r$, $\tau=-\ln\rho$ and $v(t)=\widetilde{v}(-\ln r)$, which transforms (8) into
@@
where one minimizes over all the functions such that the numerator is finite. The main theorem of [MR3694655, Theorem 1.1] states that for each $L>0$ the functional $\mathcal{E}_{s,L}$ in (11) has a nonzero minimizer. Furthermore, when $0<L\ll 1$ is sufficiently small, this minimizer is a positive constant; however, it is not constant when $L\gg 1$ is sufficiently large.
@@
$\bar{u}_{\varepsilon}(x)=\mathfrak{F}^{-1}(v_{\varepsilon})(x)=|x|^{\frac{2s-n}{2}}\bar{v}_{\varepsilon}(-\ln|x|).$
@@ 2.3. Morse index and stability
We discuss the Morse index of the standard Aubin–Talenti bubbles classified by Chen, Li and Ou [MR2131045, Theorem 1.1]. Let us recall that this family of bubbles of order $s\in(0,n/2)$ solving ($\overline{\mathcal{Q}}_{n,s}$) such that $\lim_{x\to 0}u(x)<\infty$ is given by
@@
$v_{t_{0},\varepsilon}(t)=v_{\circ}\left(\frac{t-t_{0}}{\varepsilon}\right)=\hat{c}_{n,s}\varepsilon\cosh({t-t_{0}})^{\frac{2s-n}{2}}$
for some $t_{0}\in\mathbf{R}^{n}$ and $\varepsilon>0$, where $v_{0,1}(t)=v_{\circ}(t)=\hat{c}_{n,s}\cosh(t)^{\frac{2s-n}{2}}$.
The main result of Dávila, del Pino and Sire [DDS, Theorem 1.1] characterizes the kernel of
@@
In fact, they are generated by the Jacobi fields arising from the action of two conformal transformations, namely translations and dilations. Later, Andrade, Wei and Ye [Andrade_Wei_Ye, Lemma 4.6] (see also [MR4929821, Lemma 5.1]) extended this nondegeneracy result to the full range of exponents, namely $s\in(0,\frac{n}{2})$. Changing from the Euclidean gauge to the cylindrical gauge transforms the spherical linearized operator into
@@
On a related note, we observe that the results of [Azahara_Sergio_Ruiz, Theorem 1.2] prove that all the Delaunay solutions (and its deformations) $\bar{u}_{\varepsilon}\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$ and $\bar{v}_{\varepsilon}\in\mathcal{C}^{\infty}(\mathbf{R})$ described above have infinite Morse index.
@@ 3.1. Upper bound estimate
In the case that $\Lambda$ is finite, this result follows from [MR3366748, Proposition 3.1]. This more general proof when $\Lambda$ is a closed set with zero capacity follows the model first used by Pollack in [MR1266101, Theorem 4.3] and has since been adapted in a number of other places, such as [Marques_iso_blowup], [MR2219215], [chang-han-yang], and [cmpt_6th_order]. Since this same argument has appeared now a number of times in the literature, we only provide its sketch.
@@
We argue by contradiction, adapting the blow-up argument of Pollack [MR1266101, Theorem 4.3], combined with the convexity argument of Schoen [MR1173050] and its conformal extension by Qing and Raske [MR2219215, Theorem 4.1].
@@
where $\widehat{\Lambda}$ is a finite set. Since the singular set $\widehat{\Lambda}$ has capacity zero, the results of [MR4438901] imply that $u$ is $s^{\prime}$-superharmonic for every $s^{\prime}\in(0,s)$.
Next, we define the conformal metric $g_{u}=u^{\frac{4}{n-2s}}g_{\rm flat}$. By applying the moving planes method of Schoen [MR1173050] adapted to the fractional setting by Qing and Raske [MR2219215, Theorem 4.1], we deduce that every Euclidean ball in the domain $\mathbf{R}^{n}\setminus\widehat{\Lambda}$ is strictly mean-convex with respect to $g_{u}$. This key convexity property will be used to obtain a contradiction.
@@
Because the singular sets $\widehat{\Lambda}_{k}\subset\mathbf{R}^{n}$ diverge to infinity in the rescaled variables, and thanks to the uniform a priori Harnack inequality from [TanXiong2011, Theorem 1.2], we obtain local $\mathcal{C}^{2,\alpha}$ estimates for $w_{k}$ on compact subsets. Passing to a subsequence, we conclude that $w_{k}\to w_{\infty}$ in $\mathcal{C}^{2}_{\mathrm{loc}}(\mathbf{R}^{n})$, where $w_{\infty}\in\mathcal{C}^{2}(\mathbf{R})$ satisfies $w_{\infty}(0)={2^{\frac{2s-n}{2}}}$ and solves
@@
By the classification result of Chen, Li and Ou [MR2200258, Theorem 1.1] $w_{\infty}$ must be the standard spherical solution
@@
Finally, this contradiction proves the desired a priori upper bound (13). The constant $C$ depends only on $n$ and $s$, completing the proof. ∎

### v2

@@ 1. Introduction
where $\in(0,\frac{n}{2}), s\mathbb{P}_{s}$ is the family of fractional Paneitz operators constructed by Graham and Zworski in [20] (see also [8]) and $c_{n,s}>0$ is the normalization constant defined as
@@
We are mostly interested in this equation on subdomains of the sphere, i.e. $\Omega:=\mathbf{S}^{n}\setminus\Lambda$ with $\Lambda\subset\mathbf{S}^{n}$ a closed set. We denote the usual round metric as $g_{\circ}$. In this locally conformally flat setting, $\mathbb{P}_{s}$ has the explicit form (originally given by Branson in [5]) given by
@@
Later, we will see a similar expansion for $\mathbb{P}_{s}$ in the cylindrical setting, which was first derived in [17].
@@
$\hat{c}_{n,s}=2^{\frac{2s-n}{2}}\left(\frac{\Gamma\left(\frac{1}{2}\left(\frac{n}{2}+s\right)\right)}{\Gamma\left(\frac{1}{2}\left(\frac{n}{2}-s\right)\right)}\right)^{\frac{2s-n}{2s}}\left(\frac{\Gamma\left(\frac{n}{2}-s\right)}{\Gamma\left(\frac{n}{2}+s\right)}\right)^{\frac{2s-n}{4s}}.$
Here we remark that $\hat{c}_{n,s}>1$ (cf. [17, Proposition 6.1]).
@@
In previous work, DelaTorre, del Pino, González and Wei [16] constructed a family of solutions on the twice-punctured sphere. These solutions are, after an appropriate conformal dilation, all rotationally symmetric. They generalize the well-known Delaunay metrics in the scalar curvature setting, which were found by Schoen [32], and we call them Delaunay metrics as well. A Delaunay metric is uniquely determined by its fundamental period, or equivalently by its necksize. More precisely, the fractional Delaunay family can be equivalently characterized as the set of even, positive, periodic solutions $v_{\varepsilon}:\mathbf{R}\to(0,\infty)$ of a nonlocal initial value problem (see ($\widehat{\mathcal{Q}}_{n,s}$)) that fixes the necksize parameter $\varepsilon=\min_{t\in\mathbf{R}}v_{\varepsilon}(t).$ For each necksize $\varepsilon\in(0,1]$, the Euclidean conformal factor
@@
$g_{\varepsilon}=U_{\varepsilon}^{\frac{4}{n-2s}}g_{\circ}\quad{\rm and}\quad u_{\varepsilon}=U_{\varepsilon}\cdot u_{\circ}.$
In what follows, we call $u\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$ a singular solution to ($\overline{\mathcal{Q}}_{n,s}$) if it does not extend continuously to the origin, i.e. $\lim_{x\to 0}u(x)=\infty$; otherwise, we call it a nonsingular solution.
Our first main result is a partial classification of singular solutions to the fractional Yamabe equation, which can be stated as
@@
Let $n\in\mathbf{N}$ with $n\geq 3$. There exists $\delta>0$ such that if $s\in(1-\delta,1)$ and $u\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$ is a singular solution to ($\overline{\mathcal{Q}}_{n,s}$), then there exists $\varepsilon\in(0,1]$ and $R>0$ such that
$u=\bar{u}_{\varepsilon,R}\quad{\rm with}\quad\bar{u}_{\varepsilon,R}(x)=R^{\frac{2s-n}{2}}u_{\varepsilon}\left(\frac{x}{R}\right),$
where $\bar{u}_{\varepsilon}\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$ is the Euclidean Delaunay solution of necksize $\varepsilon$ ((see (12))).
@@
Let $n\in\mathbf{N}$ with $n\geq 3$ and let $p,q\in\mathbf{S}^{n}$ with $p\neq q$. There exists $\delta>0$ such that if $s\in(1-\delta,1)$ then each element of $\mathcal{M}_{\{p,q\}}$ is a Delaunay metric.
Alternatively, we could say that $\mathcal{M}_{\{p,q\}}$ is a half-open interval in the real line, where the parameter represents the minimal period of the corresponding Delaunay solution or (equivalently) the minimum value of the Delaunay solution, i.e. the necksize.
We use the same technique as in the proof of Theorem 1 to prove that all Delaunay solutions are nondegenerate. This condition is easiest to define in the cylindrical gauge. For any $v:\mathbf{R}\times\mathbf{S}^{n-1}\rightarrow(0,\infty)$ solution to ($\widehat{\mathcal{Q}}_{n,s}$), we set the associated linearized operator as
@@
We say that $v$ is nondegenerate if $w\in L^{2}(\mathbf{R}\times\mathbf{S}^{n-1})$ and $\mathbb{L}_{v}(w)=0$ implies $w\equiv 0$.
In this setting, our second main result is
@@
Let $n\in\mathbf{N}$ with $n\geq 3$. There exists $\delta>0$ such that if $s\in(1-\delta,1)$ and $v:\mathbf{R}\times\mathbf{S}^{n-1}\rightarrow(0,\infty)$ satisfies ($\widehat{\mathcal{Q}}_{n,s}$), then $v$ is nondegenerate.
In the scalar curvature setting, the conformal factor satisfies a second order ODE, and so the uniqueness follows very quickly from phase-plane analysis. However, as demonstrated in [17], one cannot apply these techniques directly in the nonlocal setting. Nonetheless, we can prove uniqueness if $s$ is close to, but less than one.
Our proof of uniqueness relies on two new ideas. In § 3.2, we use the fact that the spherical solution has Morse index one to derive a sharp a priori upper bound for a solution to the $Q$-curvature equation of order $s$ on a twice-punctured sphere, with the spherical solution giving us a universal upper bound. This allows us to associate a Delaunay solution with the same $L^{\infty}$-norm to any solution on the twice-punctured sphere. We prove Theorem 1 by contradiction, assuming there exists a sequence $s_{k}\nearrow 1$ and solutions $v_{k}$ that are not Delaunay solutions. Letting $\overline{v}_{k}$ be the matched Delaunay solution, in § 4.1 we prove that a rescaling of the difference $v_{k}-\overline{v}_{k}$ converges locally to a solution of a homogeneous second order ODE. In § 4.2 we extend this local convergence to global convergence. This limit must be zero by the conditions of the ODE, which contradicts the normalization of our rescaling. The passage from local convergence to global convergence relies on a sharp estimate for the integral kernel and the solution operator of the nonlocal equation, derived in [16] and described below in § 2.2. We expect this strategy to be robust and to have applications in other fractional geometric problems, and indeed we adapt our technique to prove Theorem 3 in § 5. We conclude this paper with a brief outlook in § 6. First, in § 6.1, we discuss work in progress [1] in which we extend the classification to $s\in(1,1+\delta)$, where once again $\delta>0$ is sufficiently small. Next, in § 6.2, we describe some possible applications of our classification result, including refined asymptotics of solutions with isolated singularities and a general compactness result.
@@ 2. Qualitative properties of Delaunay solutions
In this section, we first describe the reduction to a one-dimensional problem. Next, we describe the Delaunay solutions. Finally, we recall that Dávila, Del Pino and Sire [15] (see also [4]) proved that the standard bubble has Morse index one.
@@ 2.1. Rotational symmetry
One can now apply [6] to see that $u$ is rotationally invariant. Alternatively, one can argue as follows. Ngô and Ye [28] prove that, because the singular set $\{0\}$ has capacity 0, the function $u$ is $s^{\prime}$-superharmonic for each $0<s^{\prime}<s$. One can now use the method of moving planes to show that $u$ is radial. One can also use the method of moving planes in integral form (see [13] and [12]) to show the radial symmetry directly.
@@ 2.2. Existence of Delaunay solutions
The construction in [16] starts with the Ansatz that the solution $u$ has the form
@@
where $\widetilde{v}:(0,\infty)\rightarrow(0,\infty)$ is a positive smooth function. In fact, one can use the a priori bounds of [6] to show that $\widetilde{v}$ is bounded, and so the integral operators written below are convergent. This change of variables transforms ($\overline{\mathcal{Q}}_{n,s}$) into
@@
is a positive constant; in fact, our normalizations imply $A=c_{n,s}$ (see [17, Proposition 2.7]). Finally, we complete the Emden-Fowler change of variables by letting $t=-\ln r$, $\tau=-\ln\rho$ and $v(t)=\widetilde{v}(-\ln r)$, which transforms (8) into
@@
where one minimizes over all the functions such that the numerator is finite. The main theorem of [16, Theorem 1.1] states that for each $L>0$ the functional $\mathcal{E}_{s,L}$ in (11) has a nonzero minimizer. Furthermore, when $0<L\ll 1$ is sufficiently small, this minimizer is a positive constant; however, it is not constant when $L\gg 1$ is sufficiently large.
@@
$\bar{u}_{\varepsilon}(x)=\mathfrak{F}^{-1}(v_{\varepsilon})(x)=|x|^{\frac{2s-n}{2}}\bar{v}_{\varepsilon}(-\ln|x|).$ (12)
@@ 2.3. Morse index and stability
We discuss the Morse index of the standard Aubin–Talenti bubbles classified by Chen, Li and Ou [11, Theorem 1.1]. Let us recall that this family of bubbles of order $s\in(0,n/2)$ solving ($\overline{\mathcal{Q}}_{n,s}$) such that $\lim_{x\to 0}u(x)<\infty$ is given by
@@
$v_{t_{0},\varepsilon}(t)=v_{\circ}\left(\frac{t-t_{0}}{\varepsilon}\right)=\hat{c}_{n,s}\cosh\left(\frac{t-t_{0}}{\varepsilon}\right)^{\frac{2s-n}{2}}$
for some $t_{0}\in\mathbf{R}$ and $\varepsilon>0$, where $v_{0,1}(t)=v_{\circ}(t)=\hat{c}_{n,s}\cosh(t)^{\frac{2s-n}{2}}$.
The main result of Dávila, del Pino and Sire [15, Theorem 1.1] characterizes the kernel of
@@
In fact, they are generated by the Jacobi fields arising from the action of two conformal transformations, namely translations and dilations. Later, Andrade, Wei and Ye [4, Lemma 4.6] (see also [10, Lemma 5.1]) extended this nondegeneracy result to the full range of exponents, namely $s\in(0,\frac{n}{2})$. Changing from the Euclidean gauge to the cylindrical gauge transforms the spherical linearized operator into
@@
On a related note, we observe that the results of [14, Theorem 1.2] prove that all the Delaunay solutions (and its deformations) $\bar{u}_{\varepsilon}\in\mathcal{C}^{\infty}(\mathbf{R}^{n}\setminus\{0\})$ and $\bar{v}_{\varepsilon}\in\mathcal{C}^{\infty}(\mathbf{R})$ described above have infinite Morse index.
@@ 3.1. Upper bound estimate
In the case that $\Lambda$ is finite, this result follows from [6, Proposition 3.1]. This more general proof when $\Lambda$ is a closed set with zero capacity follows the model first used by Pollack in [29, Theorem 4.3] and has since been adapted in a number of other places, such as [26], [30], [9], and [3]. Since this same argument has appeared now a number of times in the literature, we only provide its sketch.
@@
We argue by contradiction, adapting the blow-up argument of Pollack [29, Theorem 4.3], combined with the convexity argument of Schoen [33] and its conformal extension by Qing and Raske [30, Theorem 4.1].
@@
where $\widehat{\Lambda}$ is a finite set. Since the singular set $\widehat{\Lambda}$ has capacity zero, the results of [28] imply that $u$ is $s^{\prime}$-superharmonic for every $s^{\prime}\in(0,s)$.
Next, we define the conformal metric $g_{u}=u^{\frac{4}{n-2s}}g_{\rm flat}$. By applying the moving planes method of Schoen [33] adapted to the fractional setting by Qing and Raske [30, Theorem 4.1], we deduce that every Euclidean ball in the domain $\mathbf{R}^{n}\setminus\widehat{\Lambda}$ is strictly mean-convex with respect to $g_{u}$. This key convexity property will be used to obtain a contradiction.
@@
Because the singular sets $\widehat{\Lambda}_{k}\subset\mathbf{R}^{n}$ diverge to infinity in the rescaled variables, and thanks to the uniform a priori Harnack inequality from [35, Theorem 1.2], we obtain local $\mathcal{C}^{2,\alpha}$ estimates for $w_{k}$ on compact subsets. Passing to a subsequence, we conclude that $w_{k}\to w_{\infty}$ in $\mathcal{C}^{2}_{\mathrm{loc}}(\mathbf{R}^{n})$, where $w_{\infty}\in\mathcal{C}^{2}(\mathbf{R})$ satisfies $w_{\infty}(0)={2^{\frac{2s-n}{2}}}$ and solves
@@
By the classification result of Chen, Li and Ou [12, Theorem 1.1] $w_{\infty}$ must be the standard spherical solution
@@
Finally, this contradiction proves the desired a priori upper bound (14). The constant $C>0$ depends only on $n$ and $s$, completing the proof. ∎